# Massive deformations of supersymmetric Yang-Mills matrix models

### v1

@@ Abstract
We systematically classify all supersymmetry-preserving mass deformations of SYM matrix models in all dimensions ($D=3,4,6,10$). In $D=10$ the polarized IKKT model emerges as the only possible deformation. Moreover, in the particular case $D=4$ we identify two massive, sign-problem-free models, making them attractive candidates for non-perturbative numerical studies.
@@ 1 Introduction
Among the most revolutionary ideas in modern theoretical physics, holography suggests that gravity could emerge from a fundamental gauge theory description Maldacena:1997re ; tHooft:1993dmi ; Aharony:1999ti ; Itzhaki:1998dd ; Witten:1998qj ; Gubser:1998bc . A particularly intriguing application arises in matrix quantum mechanics models like BFSS and BMN Berenstein:2002jq ; Banks:1996vh , where gravitational phenomena are expected to emerge from a large-$N$ quantum mechanical system, avoiding the subtleties of quantum field theories.
In this work, however, we focus on an even more minimal setting where we believe gravity can emerge: super-Yang-Mills (SYM) matrix models. We define this class of matrix models by taking pure super-Yang–Mills theory in dimensions $D=3,4,6\text{ or }10$ and dimensionally reducing it all the way to a single point (zero dimension). Any deformation is allowed, provided it keeps the gauge invariance intact and preserves the full set of original supersymmetries (possibly deformed).
The most studied holographic model in this class is the IKKT model Ishibashi_1997 . Originally introduced as a non-perturbative description of type IIB string theory Ishibashi_1997 , it was later reinterpreted in the usual decoupling argument Ooguri:1998pf . Another interesting holographic model is the polarized IKKT model Bonelli_2002 ; Hartnoll:2024csr , a mass-deformed version of the IKKT model. Its advantage is to possess a tunable coupling, and infinitely many well-defined observables. It has many saddles, each of which should be dual to a specific ten-dimensional Euclidean gravitational background Komatsu:2024bop ; Komatsu:2024ydh .
Let us now discuss the structure of these matrix models. Upon dimensional reduction, the $SU(N)$ SYM gauge fields reduce to a set of $D$ Hermitian (traceless) $N\times N$ matrices $X_{I}$, one for each spacetime direction, and their supersymmetric partners reduce to a set of $\mathcal{N}=2(D-2)$ fermionic Hermitian (traceless) $N\times N$ matrices $\psi_{\alpha}$. Since the matrices have no longer any coordinate dependence, the path integral reduces to a finite number of matrix integrals. Namely, correlation functions and the partition function take the form [^1] The normalization conventions for the measure are detailed in appendix E.
@@
However, this simplicity is not without cost. Even though the models look simple and well-defined, carrying out analytical or numerical computations presents some practical difficulties. One first complication is that even though there are finitely many matrix integrals (at fixed $N$), 10 bosonic and 16 fermionic matrices still yield a large number of variables. Another complication is the so-called sign problem. In all models we consider, the action $S$ is quadratic in the fermionic variables. When considering bosonic observables, the fermionic integral can be performed exactly, resulting in a Pfaffian of a matrix $\mathcal{M}(X)$ of size $\mathcal{N}(N^{2}-1)\times\mathcal{N}(N^{2}-1)$. The expectation value of a bosonic observable $f(X)$ then takes the form
@@
where $S_{\mathrm{bos}}$ is the bosonic part of $S$. Unfortunately, in $D=10$, the Pfaffian $\mathrm{Pf}\mathcal{M}(X)$ is sign-changing. This complicates the use of numerical approaches such as Monte Carlo simulations Krauth:1998xh ; Fujii:2013sra . Furthermore, this lack of positivity means that observables such as $\mathrm{Tr}(X_{I}X_{I})^{p}$ are not necessarily positive, obstructing the use of techniques like the bootstrap Lin:2024vvg ; Lin:2020mme .
For these reasons, it is valuable to explore simplified, toy models that retain the essential features of the polarized IKKT model but are more tractable. In this work, we classify all such models. Specifically, we identify all possible mass deformations [^2] By mass-deformation, we mean any model that can be written has $S=S_{\mathrm{0}}+\frac{1}{g_{\mathrm{YM}}^{2}}\mu S_{1}+\frac{1}{g_{\mathrm{YM}}^{2}}\mu^{2}S_{2}+\dots$ where $S_{0}$ is the action of the pure SYM matrix model. of pure SYM matrix models in dimensions $D=3,4,6,10$ that preserve $SU(N)$ gauge symmetry and are invariant under $\mathcal{N}=2,4,8,16$ (deformed) supersymmetries, respectively. We restrict to models whose actions have a real bosonic part. A similar classification was performed in the context of matrix quantum mechanics Kim:2006wg , focusing on supersymmetric mass deformations of SYM quantum mechanics (including the BFSS model). [^3] Note that the models we construct are not dimensional reduction of these models. This is because naive dimensional reduction of BMN and its lower-dimensional cousins yields non-supersymmetric models in $0+0$ dimensions. Crucially, in these mass-deformed matrix quantum mechanics, the supersymmetry transformations are time-dependent. The main idea allowing this classification is discussed in section 3.
@@
Massive SYMmatrix $\begin{array}[]{c}\text{Massive SYM}\\[-2.0pt] \text{matrix models}\end{array}$ $\mathcal{N}$ $\begin{array}[]{c}\text{Residual}\\[-2.0pt] \text{symmetry}\end{array}$ Superalgebra inLorentzian $\begin{array}[]{c}\text{Superalgebra in}\\[-2.0pt] \text{Lorentzian signature}\end{array}$ $\begin{array}[]{c}\text{Deformation}\\[-2.0pt] \text{parameter}\end{array}$ $\begin{array}[]{c}\text{Pfaffian}\\[-2.0pt] \text{positivity}\end{array}$
$D=10$ 16 $SO(3)\times SO(7)$ $F(4)$ or $F^{1}(4)$ $\mu$ No
$D=6$ type I 8 $SO(3)\times SO(3)$ $D(2,1,\alpha=\frac{1}{2})$ $\mu$ No
$D=6$ type II 8 $SO(5)$ $OSp^{*}(2|4)$ or $OSp^{*}(2|2,2)$ $\mu$ No
$D=4$ type I 4 $SO(4)$ $OSp(1|2)\oplus OSp(1|2)$ $\mu$ Yes
$D=4$ type II 4 $SO(3)$ $SU(2|1)$ or $SU(1,1|1)$ $\mu_{1},\mu_{2}$ Yes
$D=3$ 2 $SO(3)$ $OSp(1|2)$ $\mu_{1},\mu_{2}$ No
@@
The parameters $\mu,\mu_{1},\mu_{2}$ are parameters of mass dimension one. The number $D$ denotes the number of bosonic matrices $X_{I}$ while $\mathcal{N}$ denotes the number of fermionic matrices $\psi_{\alpha}$, and correspondingly the number of supersymmetries. We refer to $D$ as the dimension of the matrix model, since we interpret $X^{I}$ as some $D$-dimensional target space coordinate. A mass deformation typically reduces the original $SO(D)$ global symmetry to a subgroup. The residual subgroup is indicated in the third column.
The models are written in Euclidean signature, yet the fermions satisfy Lorentzian reality conditions as we discuss in section 2. To recover a Lorentzian theory one must decide which coordinate serves as time. For instance, the Euclidean group $SO(3)\times SO(7)$ can be continued either to $SO(2,1)\times SO(7)$ or to $SO(3)\times SO(6,1)$. In the table we write the associated superalgebras as $F(4)$ or $F^{1}(4)$, respectively, and proceed analogously in the other dimensions. Our notation for the superalgebras follows VanProeyen:1999ni . The underlying simple super-Lie algebras were first classified by Kac Kac:1977em ; Kac:1977qb .
The $D=4$ models stand out as especially promising toy models. As indicated in the last column, these are the only models that do not suffer from the sign problem. Therefore, standard Monte Carlo or bootstrap methods should be applicable. It is worth noting that they remain highly rich in structure. For instance, the type II model, like the polarized IKKT model, possesses many saddle points.
We close this introduction with an important caveat: once mass-deformed, pure SYM matrix models cannot be recovered. In every mass-deformed model we have studied, the partition function blows up as the mass parameter is taken to zero, even though the undeformed partition function is known to converge. This discontinuity arises from the fermionic mass terms, a point first noted in Austing:2001ib and reviewed in (Komatsu:2024ydh, , App.B).
This paper is structured as follows. Section 2 revisits pure SYM matrix models, detailing convergence properties and summarizing a few known analytic results. Section 3 presents the idea that allows the classification of the mass-deformed models. Section 4 reviews the polarized IKKT model, outlining its saddle-point structure and some known analytic results. Sections 5-7 present the new models we construct in dimensions $D=6,4,3$ and provide a few analytic results in the low-mass and large-mass limits. The systematic classification of all models, the $D=4$ Pfaffian positivity proof, and other technical derivations are collected in the appendices.
@@ 2 Review of massless SYM matrix models
Pure SYM matrix models can exist only in $D=2,3,4,6,10$ because pure SYM does not exist in other dimensions. [^4] One argument is in terms of physical degrees of freedom (Tong:SupersymmetricFT, , Sec.2.4.3). A more technical argument is that the invariance of the action under supersymmetry relies on a $\psi^{3}$ term vanishing. This vanishes only thanks to Fierz identities that are valid in these specific dimensions. However, it is trivial in $D=2$ because the Pfaffian (2) vanishes identically.
@@
The gamma matrices $\Gamma^{I}$ satisfy the $SO(D)$ Clifford algebra. Since the Euclidean model is obtained from Wick rotating the Lorentzian model, fermions $\psi$ obey the reality conditions of $SO(D-1,1)$. For example the spinors are Majorana-Weyl in ten dimensions. [^6] Note that the Majorana condition is not invariant under Euclidean $SO(10)$ rotations, as we discuss in appendix D. However, this issue is resolved after integrating out the fermions, since the resulting Pfaffian is polynomial in $X_{I}$ and can safely be analytically continued Austing:2001ib ; Krauth:1998xh . In fact, the model would not be supersymmetric if we used $SO(10)$ Majorana fermions instead of $SO(9,1)$ Majorana-Weyl. The fermion conjugate is defined by $\bar{\psi}\equiv\psi^{\top}\mathcal{C}$, where $\mathcal{C}$ is the conjugation matrix. We refer the reader to appendix D for explicit constructions of $\mathcal{C}$, $\Gamma^{I}$ and $\psi$.
@@
Since the bosonic action $S_{\mathrm{bos}}=-\frac{1}{4}\mathrm{Tr}[X_{I},X_{J}]^{2}$ vanishes for commuting (e.g. diagonal) matrices, one may be worried that the partition function $Z$ in (2) diverges. To address this, let us first consider the purely bosonic model $S_{\mathrm{bos}}$. Even though it stretches to infinity, the region in which $e^{-S_{\mathrm{bos}}}$ stays $\mathcal{O}(1)$ can still have finite volume, provided it narrows rapidly enough at large distances. [^7] For a toy integral that mimics this feature, refer to (Komatsu:2024ydh, , Sec.2.1). This means that the bosonic integral $\int[dX_{I}]e^{-S_{\mathrm{bos}}}$ does not necessarily diverge. After the careful treatment of Austing:2001bd ; Austing:2001ib , one obtains the following convergence properties for bosonic correlators of the $SU(N)$ bosonic model in $D\geq 3$ dimensions,
@@
where in the second line we are excluding the values of $D$ and $N$ for which $Z_{\mathrm{bos.}}$ diverges. In fact, any polynomial in $X$ of degree less then the critical value given above has convergent expectation value. In the large $N$-limit, this means that all observables of finitely many matrices are well-defined.
@@
[…]
@@
The sum $\sum_{m|N}$ runs over all divisors of $N$ and the prefactor, determined in Krauth:1998xh , has been adapted to our conventions (see appendix E). In the IKKT case ($D=10$), the integral ceases to be analytic in $N$. Instead, it can be rewritten using the divisor function $\sigma_{2}(N)=\sum_{m\mid N}m^{2}$ (Komatsu:2024ydh, , Sec.2.1). The properties of $\sigma_{2}(N)$ have been analysed in the mathematical literature hardy1979introduction .
@@
$\begin{gathered}\langle\mathrm{Tr}X_{1}^{2}\rangle_{D=10}=\frac{16}{25\sqrt{\pi}}\,,\qquad\langle\mathrm{Tr}X_{1}^{4}\rangle_{D=10}=\frac{9}{40}\,,\qquad\langle\mathrm{Tr}X_{1}^{6}\rangle_{D=10}=\frac{3}{8\sqrt{\pi}}\,,\\ \langle\mathrm{Tr}X_{1}^{8}\rangle_{D=10}=\frac{297}{1024}\,,\qquad\langle\mathrm{Tr}X_{1}^{10}\rangle_{D=10}=\frac{1089}{1024\sqrt{\pi}}\,,\qquad\langle\mathrm{Tr}X_{1}^{12}\rangle_{D=10}=\frac{184041}{81920}\,,\end{gathered}$ (14)

### v2

@@ Abstract
We systematically classify all supersymmetry-preserving mass deformations of SYM matrix models in all dimensions ($D=3,4,6,10$). In $D=10$, the polarized IKKT model emerges as the only possible deformation. In $D=4$, we identify two massive models without a sign problem, making them attractive candidates for non-perturbative numerical studies.
@@ 1 Introduction
Among the most revolutionary ideas in modern theoretical physics, holography suggests that gravity could emerge from a fundamental gauge theory description Maldacena:1997re; tHooft:1993dmi; Aharony:1999ti; Itzhaki:1998dd; Witten:1998qj; Gubser:1998bc. A particularly intriguing application arises in matrix quantum-mechanical models like BFSS and BMN Berenstein:2002jq; Banks:1996vh, where gravitational phenomena are expected to emerge from a large-$N$ quantum mechanical system, avoiding the subtleties of quantum field theories.
In this work, however, we focus on an even more minimal setting where we believe gravity can emerge: super-Yang-Mills (SYM) matrix models. We define this class of matrix models by taking pure super-Yang–Mills theory in dimensions $D=3,4,6\text{ or }10$ and dimensionally reducing it all the way to a single point (zero dimension). We then consider deformations that preserve $SU(N)$ gauge invariance and the full set of original supersymmetries (possibly in a deformed form).
The most studied holographic model in this class is the IKKT model Ishibashi_1997. Originally introduced as a non-perturbative description of type IIB string theory Ishibashi_1997, it was later reinterpreted in terms of the usual decoupling limit Ooguri:1998pf. Another interesting holographic model is the polarized IKKT model Bonelli_2002; Hartnoll:2024csr, a mass-deformed version of the IKKT model. A key advantage is that it possesses a tunable coupling, and infinitely many well-defined observables. It has many saddles, each of which is expected to be dual to a specific ten-dimensional Euclidean gravitational background Komatsu:2024bop; Komatsu:2024ydh.
Let us now discuss the structure of these matrix models. Upon dimensional reduction, the $SU(N)$ SYM gauge fields reduce to a set of $D$ Hermitian (traceless) $N\times N$ matrices $X_{I}$, one for each spacetime direction, and their supersymmetric partners reduce to a set of $\mathcal{N}=2(D-2)$ fermionic Hermitian (traceless) $N\times N$ matrices $\psi_{\alpha}$. Since the matrices no longer have any coordinate dependence, the path integral reduces to a finite number of matrix integrals. Namely, correlation functions and the partition function take the form [^1] The normalization conventions for the measure are detailed in appendix E.
@@
However, this simplicity is not without cost. Even though the models look simple and well-defined, carrying out analytical or numerical computations presents some practical difficulties. A first complication is that, even though there are finitely many matrix integrals (at fixed $N$), 10 bosonic and 16 fermionic matrices still yield a large number of variables. Another complication is the so-called sign problem. In all models we consider, the action $S$ is quadratic in the fermionic variables. When considering bosonic observables, the fermionic integral can be performed exactly, resulting in the Pfaffian of an $\mathcal{N}(N^{2}-1)\times\mathcal{N}(N^{2}-1)$ matrix $\mathcal{M}(X)$. The expectation value of a bosonic observable $f(X)$ then takes the form
@@
where $S_{\mathrm{bos}}$ is the bosonic part of $S$. Unfortunately, in $D=10$, the Pfaffian $\mathrm{Pf}\mathcal{M}(X)$ is sign-changing. This complicates the use of numerical approaches such as Monte Carlo simulations Krauth:1998xh; Fujii:2013sra. [^2] It is worth noting, however, that promising results have been achieved in this direction through approaches such as the complex Langevin method and the Lefschetz thimble technique. Fujii:2013sra; Anagnostopoulos:2020xai; Anagnostopoulos:2017gos; Anagnostopoulos:2022dak; Chou:2025moy Furthermore, this lack of positivity means that observables such as $\mathrm{Tr}(X_{I}X_{I})^{p}$ are not necessarily positive, obstructing the use of techniques like the bootstrap Lin:2024vvg; Lin:2020mme.
For these reasons, it is valuable to explore simplified toy models that retain the essential features of the polarized IKKT model but are more tractable. In this work, we classify all such models. Specifically, we identify all possible mass deformations [^3] By mass deformation, we mean any model that can be written as $S=S_{\mathrm{0}}+\frac{1}{g_{\mathrm{YM}}^{2}}\mu S_{1}+\frac{1}{g_{\mathrm{YM}}^{2}}\mu^{2}S_{2}+\dots$ where $S_{0}$ is the action of the pure SYM matrix model. of pure SYM matrix models in dimensions $D=3,4,6,10$ that preserve $SU(N)$ gauge symmetry and are invariant under $\mathcal{N}=2,4,8,16$ (deformed) supersymmetries, respectively. We restrict to models whose actions have a real bosonic part. A similar classification was performed in the context of matrix quantum mechanics Kim:2006wg, focusing on supersymmetric mass deformations of SYM quantum mechanics (including the BFSS model). [^4] Note that the models we construct are not dimensional reductions of these models. Naive dimensional reduction of BMN and its lower-dimensional cousins would yield non-supersymmetric models in $0+0$ dimensions. Crucially, in these mass-deformed matrix quantum mechanics, the supersymmetry transformations are time-dependent.
@@
$\begin{array}[]{c}\text{Massive SYM}\\[-2.0pt] \text{matrix models}\end{array}$ $\mathcal{N}$ $\begin{array}[]{c}\text{Residual}\\[-2.0pt] \text{symmetry}\end{array}$ Superalgebra $\begin{array}[]{c}\text{Deformation}\\[-2.0pt] \text{parameter}\end{array}$ $\begin{array}[]{c}\text{Pfaffian}\\[-2.0pt] \text{positivity}\end{array}$ [^5] The Pfaffian positivity is discussed in Euclidean signature.
$D=10$ 16 $\begin{array}[]{c}SO(2,1)\times SO(7)\\[-2.0pt] SO(3)\times SO(6,1)\end{array}$ $\begin{array}[]{c}F(4)\\[-2.0pt] F^{1}(4)\end{array}$ $\mu$ No
$D=6$ type I 8 $SO(2,1)\times SO(3)$ $D(2,1,\alpha=\frac{1}{2})$ $\mu$ No
$D=6$ type II 8 $\begin{array}[]{c}SO(5)\\[-2.0pt] SO(4,1)\end{array}$ $\begin{array}[]{c}OSp^{*}(2|4)\\[-2.0pt] OSp^{*}(2|2,2)\end{array}$ $\mu$ No
$D=4$ type I 4 $SO(3,1)$ $OSp(1|2)\oplus OSp(1|2)$ $\mu$ Yes
$D=4$ type II 4 $\begin{array}[]{c}SO(3)\\[-2.0pt] SO(2,1)\end{array}$ $\begin{array}[]{c}SU(2|1)\\[-2.0pt] SU(1,1|1)\end{array}$ $\mu_{1},\mu_{2}$ Yes
$D=3$ 2 $SO(2,1)$ $OSp(1|2)$ $\mu_{1},\mu_{2}$ No
@@
The quantities $\mu,\mu_{1},\mu_{2}$ are deformation parameters of mass dimension one. The number $D$ denotes the number of bosonic matrices $X_{I}$ while $\mathcal{N}$ denotes the number of fermionic matrices $\psi_{\alpha}$, and correspondingly the number of supersymmetries. We refer to $D$ as the dimension of the matrix model, since we interpret $X^{I}$ as $D$-dimensional target space coordinates. Although we will soon work in Euclidean signature, the fermions satisfy Lorentzian reality conditions, as discussed in section 2. For this reason, the models are labeled by their Lorentzian symmetry groups. A mass deformation typically reduces the original SO(D1,1)SO(D-1,1) Lorentz symmetry to a subgroup. The residual subgroup is indicated in the third column. Our notation for the superalgebras in the fourth column follows VanProeyen:1999ni. The underlying simple super-Lie algebras were first classified by Kac Kac:1977em; Kac:1977qb.
The $D=4$ models stand out as especially promising toy models. As indicated in the last column, these are the only models that do not suffer from the sign problem. Therefore, standard Monte Carlo or bootstrap methods should be applicable. It is worth noting that they still have a rich structure. For instance, the type II model, like the polarized IKKT model, possesses many saddle points.
There is an important caveat: once mass-deformed, pure SYM matrix models cannot be recovered. In every mass-deformed model we have studied, the partition function blows up as the mass parameter is taken to zero, even though the undeformed partition function is known to converge. This discontinuity arises from the fermionic mass terms, a point first noted in Austing:2001ib and reviewed in (Komatsu:2024ydh, App.B).
Finally, we note that a particular subset of the models constructed in this paper can be derived from SYM theories defined on the $n$-sphere $S^{n}$, by extending the results of Blau:2000xg to the case n=0n=0.666This was first noted in the case of massive SYM quantum mechanics with $n=1$ Bobev:2024gqg. In this identification, the mass parameter plays the role of the inverse of the sphere radius. This identification will be analyzed case by case in the main text.
This paper is structured as follows. Section 2 revisits pure SYM matrix models, detailing convergence properties and summarizing a few known analytic results. Section 3 presents the main idea underlying the classification of the mass-deformed models. Section 4 reviews the polarized IKKT model, outlining its saddle-point structure and some known analytic results. Sections 5-7 present the new models we construct in dimensions $D=6,4,3$ and provide a few analytic results in the low-mass and large-mass limits. The systematic classification of all models, the $D=4$ Pfaffian positivity proof, and other technical derivations are collected in the appendices.
@@ 2 Review of massless SYM matrix models
Pure SYM matrix models can exist only in $D=2,3,4,6,10$ because pure SYM does not exist in other dimensions. [^7] One argument is in terms of physical degrees of freedom (Tong:SupersymmetricFT, Sec.2.4.3). A more technical argument is that the invariance of the action under supersymmetry relies on a $\psi^{3}$ term vanishing. This vanishes only thanks to Fierz identities that are valid in these specific dimensions. However, it is trivial in $D=2$ because the Pfaffian (2) vanishes identically.
@@
The gamma matrices $\Gamma^{I}$ satisfy the $SO(D)$ Clifford algebra. Since the Euclidean model is obtained from Wick rotating the Lorentzian model, fermions $\psi$ obey the reality conditions of SO(D1,1)SO(D-1,1). For example, the spinors are Majorana-Weyl in ten dimensions. [^9] Note that the Majorana condition is not invariant under Euclidean $SO(10)$ rotations, as we discuss in appendix D. However, this issue is resolved after integrating out the fermions, since the resulting Pfaffian is polynomial in $X_{I}$ and can safely be analytically continued Austing:2001ib; Krauth:1998xh. In fact, the model would not be supersymmetric if we used $SO(10)$ Majorana fermions instead of $SO(9,1)$ Majorana-Weyl. The fermion conjugate is defined by $\bar{\psi}\equiv\psi^{\top}\mathcal{C}$, where $\mathcal{C}$ is the conjugation matrix. We refer the reader to appendix D for explicit constructions of $\mathcal{C}$, $\Gamma^{I}$ and $\psi$.
@@
Since the bosonic action $S_{\mathrm{bos}}=-\frac{1}{4}\mathrm{Tr}[X_{I},X_{J}]^{2}$ vanishes for commuting (e.g. diagonal) matrices, one may be worried that the partition function $Z$ in (2) diverges. To address this, let us first consider the purely bosonic model $S_{\mathrm{bos}}$. Even though it stretches to infinity, the region in which $e^{-S_{\mathrm{bos}}}$ stays $\mathcal{O}(1)$ can still have finite volume, provided it narrows rapidly enough at large distances. [^10] For a toy integral that mimics this feature, refer to (Komatsu:2024ydh, Sec.2.1). This means that the bosonic integral $\int[dX_{I}]e^{-S_{\mathrm{bos}}}$ does not necessarily diverge. After the careful treatment of Austing:2001bd; Austing:2001ib, one obtains the following convergence properties for bosonic correlators of the $SU(N)$ bosonic model in $D\geq 3$ dimensions,
@@
where in the second line we are excluding the values of $D$ and $N$ for which $Z_{\mathrm{bos.}}$ diverges. In fact, any polynomial in $X$ of degree less than the critical value given above has convergent expectation value. In the large $N$-limit, this means that all observables of finitely many matrices are well-defined.
@@
for any $N\geq 2$. [^11] Note that here the implication is only one way. It is not proven that all other values of $k$ or $D$ lead to divergences. In particular, the case $D=3$ is subtle. For even $N$, it turns out that the partition function trivially vanishes by examining the Pfaffian. It is conjectured that it vanishes not only for even $N$ but for all values of $N$ Krauth_1998. Note that contrary to the bosonic case, not all observables are necessarily convergent in the large $N$-limit. Fermions effectively enlarge the valleys at infinity. This is very similar to the case of BFSS, where the bosonic BFSS model has discrete spectrum Simon:1983jy, but it becomes continuous once adding fermions deWit:1988xki. [^12] We are grateful to J. Penedones for pointing out this nice analogy.
@@
The sum $\sum_{m|N}$ runs over all divisors of $N$ and the prefactor, determined in Krauth:1998xh, has been adapted to our conventions (see appendix E). In the IKKT case ($D=10$), the integral ceases to be analytic in $N$. Instead, it can be rewritten using the divisor function $\sigma_{2}(N)=\sum_{m\mid N}m^{2}$ (Komatsu:2024ydh, Sec.2.1). The properties of $\sigma_{2}(N)$ have been analyzed in the mathematical literature hardy1979introduction.
@@
[…]